How do you find the slope of 4y= 3(y+2)
Answer by Maths68(1474)

4y=3(y+2)
4y=3y+6
4y-3y=6
y=6
y=(0)x+6
Compare above equation with standard equation of the line
y=mx+b
Slope = m = 0, y-intercept = b = 6
Slope of the line is zero, it is a horizontal line.